1. An infinite non-conducting sheet has a surface charge density šœŽ = 0.10šœ‡š¶/š‘š2 on oneĀ 

side. How far apart are equipotential surfaces whose potentials differ by 50Volts?

2. As a space shuttle moves through the dilute ionized gas of Earthā€™s ionosphere, the shuttleā€™sĀ 

potentials is typically changed by -1.0V during one revolution. Assuming the shuttle is aĀ 

sphere of radius 10m, estimate the amount of charge to collects.

3. What are the charge and the charge density on the surface of a conducting sphere ofĀ 

radius 0.15m whose potential is 200V with V=0 at infinity?

4. A spherical drop of water carrying a charge of 30pC has a potential of 500V at its surfaceĀ 

(with V=0 at infinity). (a) What is the radius of the drop? (b) If two such drops of the sameĀ 

charge and radius combine to form a single spherical drop. What is the potential at theĀ 

surface of the new drop?



1
Expert's answer
2021-11-11T17:07:32-0500

1)

"E=\frac{\sigma}{2\varepsilon_0}," "\varphi=Ed,\implies"

"d=\frac{2\varphi\epsilon_0}{\sigma}=8.85~mm."

2)

"q=\frac{\varphi r}k=-1.1\cdot 10^{-9}~C."

3)

"q=\frac{\varphi r}k=-3.3\cdot 10^{-9}~C,"

"\sigma=\frac q{4\pi r^2}=-11.8\cdot 10^{-9}~\frac C{m^2}."

4)

a)

"r=\frac{kq}{\varphi}=0.54~mm,"

b)

"\varphi'=\frac{2kq}{\sqrt[3]2 r}=\sqrt[3]4\varphi=794~V."